OpenStudy (anonymous):

Solve the problem. Be sure to show all of your work. The area of a triangle with base b and height h is given by the formula A=1/2bh Find the area of a triangle with base 3 meters and height 10 meters

OpenStudy (anonymous):

A=1/2(3)(10) A=1/2(30) A=30/2=15m^2 You have to put the meters squared because it's the units

OpenStudy (amistre64):

really? they give you the formula, tell you what the variables equal and you still want someone to do it for you for free ? lol.... ok

OpenStudy (anonymous):

Hey, I just love me some math :) They're like math coco puffs :) Or lucky math charms :) Go great with milk :D

OpenStudy (anonymous):

@ amisre64 i didnt get how to do the problem and @ Mknight94 thanks!

OpenStudy (anonymous):

Welcome! Anytime you need a problem solved, I gots yo back! (btw, I'm white, so the gots yo back part, oh yeah... That right :D)

OpenStudy (anonymous):

Mknight94 lol math is the weakest for me, english is my subject! ! can u help me with the rest! please!

OpenStudy (amistre64):

...... apparently so, but its hard to fathom for me :)

OpenStudy (anonymous):

uuh, sure! :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Just shoot me up on chat, and I"ll answer em for ya

OpenStudy (anonymous):

k

OpenStudy (anonymous):

a=1/60